This is a problem from a college algebra project, and we are stuck.
A motorist drove 150 miles on country roads before driving 50 miles on mountain roads. The rate of speed on the country roads was three times the rate on mountain roads. The time spent travelling the 200 miles was 5 hours. Find the rate of the motorist on country roads.

Answer by scott8148(5891)

r*t=d __ t=d/r __ the total time is the sum of country time and mountain time

let x=country rate

[150/x]+[50/(x/3)]=5 __ multiplying by x __ 150+150=5x __ dividing by 5 __ 60=x